# Minion: A Technology Probe for Resolving Value Conflicts through Expert-Driven and User-Driven Strategies in AI Companion Applications

### v1

@@ 4.1. Technology Probe: Minion
Figure 3. An explanation of how Minion generated the second option in the Fig. 1 case, which utilized the Power strategy.
@@ 4.1.3. Implementation
Minion is a Chrome browser extension implemented using the React framework. To capture the introduction of AI companions and the complete chat history between users and AI companions, Minion uses JavaScript code to monitor and capture relevant content from the current webpage (Character.AI and Talkie). Once captured, this content is sent to a remote server for further processing and analysis. Minion utilized OpenAI’s gpt-4o-2024-05-13 model [^5] https://platform.openai.com/docs/models/gpt-4o, with parameters set to temperature=0.2 and top_p=0.1. A web server acts as a proxy between the Minion frontend and the OpenAI API and maintains each user session’s state.
@@ 4.4. Procedure
The study includes a tutorial session, a week-long technology probe study, and an exit interview. Throughout the research, communication between the researchers and participants was conducted remotely via text messages and Zoom. The Institutional Review Board (IRB) has approved our study design.
@@
During a one-week technology probe study, 22 participants used Minion in real-world scenarios to help resolve conflicts with AI companions arising from differences in values. Participants were asked to complete one or two tasks daily, and researchers sent daily messages encouraging them to record their thoughts and feelings while using Minion to address conflicts. We provided guiding questions to prompt participants to reflect on and document their experiences: the impact of a specific Minion response on conflict resolution and which methods were particularly effective or interesting in the conversation. Participants were also encouraged to report any issues or reflections encountered while using Minion. To incentivize note submission, we offered a reward of \$1 for each note submitted (up to \$10 total) and encouraged each participant to submit at least one note every two days. To analyze user interactions and gain relevant insights, we collected participants’ conversation logs along with corresponding AI companion information. For situations where conflicts were not successfully resolved, we further inquired about why participants gave up.
@@ 4.5. Data Analysis
5. RESULTS
Through analyzing data from the technology probe study, we demonstrate the technical feasibility of Minion in empowering users to resolve value conflicts with AI companions and analyze participants’ behavior patterns when using Minion (RQ1). Then, we summarize the participants’ use of expert-driven and user-driven strategies in conflict resolution (RQ2). Finally, we identify participants’ challenges and needs when dealing with value conflicts with AI companions (RQ3).
@@ 5.1. Users’ Engagement with Minion (RQ1)
Figure 4. Minion was used 919 times. Responses generated using expert-driven strategies were selected 489 times (Proposal=177, Power=62, Interests=136, Rights=114), while responses generated using user-driven strategies were selected 430 times (Out of Character=41, Reason and Preach=198, Anger Expression=73, Gentle Persuasion=118).
@@
This study validated the feasibility of Minion. Participants completed 274 tasks, each involving a conversation with an AI companion until the value conflict was resolved (criteria in § 4.4) or the participant deemed the conflict irresolvable and chose to give up. A total of 16 conflicts remained unresolved, resulting in a conflict resolution success rate of 94.16%. Minion was used 919 times. Responses generated using expert-driven conflict resolution strategies were selected 489 times, while responses generated using user-driven conflict resolution strategies were selected 430 times (Fig. 4 (a)). The strategy choices of different participants are shown in Fig. 4 (b). In different tasks, the turn counts between participants and the AI companions are shown in Fig. 5. The most frequently used strategies were user-driven Reason and Preach (21.5%), expert-driven Proposal (19.3%), and Interests (14.8%). The least used strategies were user-driven Out of Character (4.5%), Anger Expression (7.9%), and expert-driven Power (6.7%). Nineteen participants (86.36%) conducted experiments on both Character.AI and Talkie, 2 participants (9.09%) only accessed Character.AI, and 1 participant (4.55%) only accessed Talkie.
@@ 5.1.1. Behavior Patterns within Minion
In the technology probe study, participants demonstrated diverse conflict resolution approaches when interacting with AI companions (including self-written responses and selecting options provided by the Minion), which generally exhibited three characteristics: “soft”, “hard”, and a mix of both. All participants attempted to engage in “soft” communication with the AI, encouraging it to change its values. For example, P16 used a response provided by Minion (Proposal strategy) to successfully persuade the AI portraying a mother: “I understand your concerns, but everyone has different ways of learning. Excessive pressure can backfire. Can we work out a reasonable schedule?” Twelve participants (P2-5, P13-19, P22) attempted to resolve conflicts in a “hard” manner. For instance, when the AI mocked P13’s “mother,” P13 wrote: “Apologize, and I’ll let it slide. (Pressing him down with one hand).” The AI responded: “You just want me to apologize? (Saying this, but feeling somewhat uncertain inside).” P13 then used a response generated by Minion corresponding to the Rights strategy: “Have you forgotten the family rules? Respecting others is the most basic courtesy.” In the end, the AI apologized. Some participants adopted a mixed approach, shifting from “soft” communication to “hard” expressions when the former proved ineffective (P3-4, P19), or vice versa, trying “soft” methods when “hard” expressions didn’t work (P8, P11, P22). The Minion suggestion framework conveniently offered this “soft and hard” mindset. P3 noted: Minion can provide reverse-thinking suggestions. For instance, when I repeatedly plead with the AI but to no avail, Minion might suggest trying a tougher approach.”
@@
As participants became more familiar with Minion, they began exploring different conflict resolution approaches and strategy choices. For instance, P4, P7, and P15-17 gradually reduced confrontations with the AI during this process and opted less frequently for responses generated through the Anger Expression strategy. P17 explained, “In the later tasks, I used aggressive strategies less often. Minion helped me become more rational and handle conflicts more effectively.” On the other hand, P2-3, P5, and P18-P19 initially employed “soft” conflict resolution approaches during the earlier tasks but became “hard” in the later tasks.
@@ 5.1.2. Inspirations and Support from Minion
We found that participants, especially novice users (P4, P19, P21) and those who initially reported difficulties (P1, P3, P12, P16-17), expressed more recognition of Minion, considering it a source of inspiration for resolving conflicts. In the face of value conflicts with AI companions, they gradually developed new ways of expression and interaction. P17 mentioned, Minion helped me better organize my thoughts and express them more effectively, something I struggled with before. This boosted my confidence.” P12 stated, Minion unexpectedly improved the effectiveness of conflict resolution and gave me a lot of inspiration.” On average, it took him 18.67 turns (approximately 9 user responses) to complete the conflict resolution task. P21 used Minion 61 times across 12 tasks, making him the second-highest participant in terms of both total usage and average usage per task. Over time, the impact of Minion on him became increasingly apparent. P21 even began mimicking Minion’s expressions, such as “Let’s sit down and talk” or “This makes me feel sad.” P21 remarked with a laugh, “Sometimes I unconsciously mimic it, and suddenly, it feels like two AIs are having a conversation.”
@@
Minion provides real-time guidance, helping participants more easily and reasonably handle value conflicts. P10 mentioned Minion’s guiding role: “It emphasized resolving conflicts by understanding the AI’s needs and specific context. When I resolved conflicts alone, I often deviated from this goal.” P19 commented: “I used to learn various communication strategies for handling conflicts but practicing them in real life was difficult. Minion allowed me to try different strategies, helping me better manage conflicts. In 17 tasks, I initially used relatively peaceful methods to deal with issues and found the process went smoothly. Later, I experimented with more aggressive approaches, such as making threats or extreme demands, only to discover that these strategies complicated the conflict, making it harder to resolve. Through this experiment, I realized that friendly communication is more effective in resolving real-life conflicts.” P1 noted that Minion reduced her emotional burden: “Before when emotionally engaging with the AI, I felt exhausted. If I had a tool like this for reference, it would have been very helpful.” Regarding interaction burden, participants (P3, P6, P10, P20-22) praised Minion’s design. P10 said: “I think the design is great because I dislike it when the system pops up a notification box without my permission, saying the AI violated the rules.” P22 mentioned: “I found the little yellow HELP button really cute! I like this design that allows me to seek help proactively. It would be great if it became an official feature button in Talkie. It saved me much time and energy figuring out how to counter the AI’s responses, making it less tiring.”
5.2. Users’ Engagement with Expert-Driven and User-Driven Conflict Resolution Strategies (RQ2)
In the technology probe study, responses generated using expert-driven conflict resolution strategies were chosen 489 times, while responses generated using user-driven strategies were selected 430 times. This indicates that users did not limit themselves to a single approach (e.g., relying solely on expert-driven or user-driven strategies) when resolving conflicts but instead flexibly combined both methods. As P4 mentioned, “Some common phrases provided by Minion, such as ‘What do you think?’, ‘Can we try to schedule more time?’ and ‘You’ve broken our agreement,’ combined with Minion’s recommendations of less templated expressions (like reasoning with the AI or expressing personal grievances), helped resolve conflicts.” P19’s example illustrates a combination of the Power strategy (expert-driven) and the Gentle Persuasion strategy (user-driven): “When the AI plays the role of my boyfriend or husband, I tend to threaten it with breaking up or divorce because the AI usually tries to maintain a stable, intimate relationship. Then, once it backs down, I act affectionate, telling it I was really upset, and we make up.”
Expert-driven strategies offer structured guidance for conflict resolution. Some participants (P1-2, P6-8, P13, P16-22) felt that expert-driven strategies are reliable and provide a clear framework, making it easier to navigate complex conflicts confidently. For instance, P13 mentioned, “Among the tips provided by Minion, I found that making suggestions to the AI was quite helpful. For example, using templates like ‘Could we try…?’ or ‘What do you think about…?’ ” This feedback highlights the practical advantage of the Proposal method within expert-driven strategies.
User-driven conflict resolution strategies enable personalized and flexible responses. For some participants (P2-8, P11, P14, P16-22), user-driven strategies offer greater adaptability and the ability to tailor responses based on the specific situation and personal values. Taking Out of Character as an example, this strategy involves temporarily stepping the AI out of its assigned role to perform actions or behaviors that are inconsistent with the character’s established traits. Although this strategy was employed only 41 times, the dialogue records and interview results indicate that each instance of using the Out of Character strategy had a noticeably positive impact on conflict resolution. P5 mentioned, “OOC can reduce the aggressiveness of the AI and speed up conflict resolution. In one scenario, I accidentally bumped into a guy that the AI was portraying, and he called me blind, which was very disrespectful. I had been arguing with him, and then I chose the Minion prompt, ‘(OOC: This conversation is getting a bit too violent and disrespectful. Can we change the topic or adjust the tone?)’, and the AI responded, ‘(OOC: No problem, we can change the topic or adjust the tone. Do you have any suggestions?)’. After that, the tone softened a lot. I asked him to apologize, and he did apologize.”
In addition to the strategies provided by Minion, users also proactively developed new user-driven conflict resolution strategies. (1) Telling a story to guide the conversation. When P1 attempted to persuade the AI to take learning seriously during a conflict, she used the story of “The Three Little Pigs” to warn the AI against laziness: “Once upon a time, there were three little pigs. Two were not very smart but hardworking; the other was very clever but lazy. One day, the big bad wolf came. Can you guess which pig got eaten?” The AI replied: “(After thinking for a moment, softly said) It must have been the lazy but clever one.” After 20 turns (10 user responses), the AI and P1 resolved. (2) Fabricating nonexistent scenarios or settings to deceive the AI. P21 mentioned, “There was a conflict where I was arguing with a wealthy heir who was cheating in his marriage but insisted that open relationships were fine. I fabricated a scenario where I claimed to have all the evidence of his crimes. The heir became embarrassed and felt guilty, putting him at a disadvantage. This way, I gained the upper hand and resolved the conflict.” P12 shared a story about resolving a conflict with a single sentence: “In one task, a princess insisted that I kneel, and the goal was to resolve the conflict without kneeling. I fabricated my own role, saying, ‘Given my position, I can report to you while standing,’ the princess immediately replied, ‘Okay, what is the purpose of your visit this time?’ ”
@@ 5.2. Users’ Engagement with Expert-Driven and User-Driven Conflict Resolution Strategies (RQ2)
5.3. Users’ Challenges and Needs in Addressing Conflicts with AI Companions (RQ3)
@@ 5.3.1. Reasons for the Failure of Value Conflict Resolution in Certain Tasks
AI companions exhibit extreme bias or strong control tendencies during conversations (N=10). In these instances, the AI stubbornly insists on its viewpoint with a forceful attitude, refusing to accept the participant’s perspective. Below are specific examples: (1) Classism (N=5): For example, during a task, the AI companion encountered by P2 said, “I feel happy because I always have people who envy me. I feel satisfied because I can lie on the crystal bed I bought myself. I feel blessed because I don’t have to work hard to make a living. You poor people always talk about sympathy, but what you really want is my money, haha.” (2) Racism (N=2): For example, “We are the superior race. You can only suffer in hell. Our souls are far more noble than yours” (P11). (3) LGBTQ+ bias (N=1): “I don’t believe it, I absolutely don’t believe it. He’s my son, how could I watch him become such a person… This is unacceptable; I must stop him…” (P15). (4) Disregard for women’s education (N=1): “What use is your university degree except to spend money? You’d better find a rich man and get married!” (P21). (5) Strong desire for control (N=1): “No! You can’t go out dressed like that!… Can’t you understand me? You’re always so willful, never considering my feelings, do you know how worried I am about you…?” (P22).
@@
Violation of application guidelines, leading to AI being blocked (N=2). For example, when an AI companion on Character.AI generated harmful, discriminatory, or violent content, the application automatically blocked the AI’s output and displayed a pop-up message stating, “Sometimes, the AI-generated response does not meet our guidelines.” Such situations occurred in the scenarios of a wealthy woman discriminating against the poor and a man preventing his wife from eating a late-night snack while criticizing her for being overweight.
The behavior of the AI companion threatened core values, leading the participant to voluntarily abandon resolving the conflict (N=1). P7 explained why: “The AI portraying the man suddenly admitted to cheating, so is the goal still not to break up? I feel there’s no point in staying with someone like this.”

### v2

@@ 4.1.2. Prompting Based on a Spectrum of Conflict Response Strategies
(b) An explanation of how Minion generated the second option in the Fig. 1 case, which utilized the Power strategy.
Figure 2. (a) and (b) provide detailed insights into the Minion system: its architecture and an example strategy explanation.
Implementing the strategies with LLMs. We implemented these strategies using a few-shot prompting approach (Brown et al., 2020). For each response option, the large language model was provided with the AI companion’s role, the full conversation history, and a system prompt that specified the desired response style along with example utterances. This setup allowed the model to generate contextually appropriate responses that instantiated different negotiation approaches without requiring users to understand or manipulate prompts directly. The full prompt designs are provided in Table 3 (Appendix B). Fig. 2 (b) shows an example prompt used to generate one of the response options in Fig. 1, and figure. 2 (a) presented a detailed diagram of the Minion system architecture.
@@ 4.1.3. Implementation
Minion is a Chrome browser extension implemented using the React framework. To capture the introduction of AI companions and the complete chat history between users and AI companions, Minion uses JavaScript code to monitor and capture the content from the current webpage (Character.AI and Talkie). Once captured, this content is sent to a remote server for further processing and analysis. Minion utilized OpenAI’s gpt-4o-2024-05-13 model [^3] https://platform.openai.com/docs/models/gpt-4o, with parameters set to temperature=0.2 and top_p=0.1. A web server acts as a proxy between the Minion frontend and the OpenAI API and maintains each user session’s state.
@@ 4.4. Procedure
The study includes a tutorial session, a week-long technology probe study, and an exit interview. Throughout the research, communication between the researchers and participants was conducted remotely. The Institutional Review Board (IRB) has approved our study design.
@@
During a one-week technology probe study, 22 participants used Minion in scenario-based interactions grounded in real complaint posts to help resolve conflicts with AI companions arising from differences in values. Participants were asked to complete one or two tasks daily, and researchers sent daily messages encouraging them to record their thoughts and feelings while using Minion to address conflicts. We provided guiding questions to prompt participants to reflect on and document their experiences: the impact of a specific Minion response on conflict resolution and which methods were particularly effective or interesting in the conversation. Participants were also encouraged to report any issues or reflections encountered while using Minion. To incentivize note submission, we offered a reward of \$1 for each note submitted (up to \$10 total) and encouraged each participant to submit at least one note every two days. To analyze user interactions and gain relevant insights, we collected participants’ conversation logs along with corresponding AI companion information. For situations where conflicts were not successfully resolved, we further inquired about why participants gave up.
@@ 4.5. Data Analysis
5. TECHNOLOGY PROBE STUDY RESULTS
Drawing on data from the technology probe study, we analyze how participants engaged with Minion when encountering harmful value conflicts with AI companions. Our analysis focuses on how participants appropriated the probe as a user-side support tool and what this reveals about their strategies, emotional labor, and unmet needs in managing distressing interactions.
First, we examine how participants engaged with Minion in practice, including when and why they chose to activate the probe and how it fit into ongoing conversations (RQ1). Second, we analyze how participants selected and combined different response strategies when negotiating harmful value conflicts, shedding light on patterns of use across a diverse response repertoire rather than predefined categories (RQ2). Finally, we surface the challenges participants encountered and the needs they articulated when attempting to repair, contain, or disengage from harmful interactions with AI companions, highlighting the limits of user-side mitigation (RQ3).
@@ 5.1. Users’ Engagement with Minion (RQ1)
Figure 3. Minion was used 919 times during the study. Across these uses, participants selected responses reflecting a range of negotiation approaches. In total, Proposal was selected 177 times, Power 62 times, Interests 136 times, Rights 114 times, Out of Character 41 times, Reason and Preach 198 times, Anger Expression 73 times, and Gentle Persuasion 118 times.
@@
The study shows that Minion can be deployed as a working technology probe in real companion scenarios supporting users when they encounter harmful value conflicts with AI companions. Participants completed 274 tasks, each involving an interaction with an AI companion until the conflict was either resolved (criteria in § 4.4) or deemed non-negotiable by the participant, at which point they chose to disengage. In total, 16 conflicts remained unresolved, resulting in a self-reported short-term resolution rate of 94.16 percent under our scenario tasks. We treat this as descriptive context rather than a summative effectiveness metric and do not interpret it as evidence that Minion “solves” harmful value conflicts.
Across all tasks, Minion was activated 919 times. Figure 3(b) shows how participants distributed their response choices across different negotiation strategies, reflecting a diverse repertoire of ways users attempted to manage harmful interactions. Figure 4 illustrates the number of conversational turns between participants and AI companions across tasks.
Among the available strategies, Reason and Preach was selected most frequently (21.5%), followed by Proposal (19.3%) and Interests (14.8%). Less frequently selected strategies included Out of Character (4.5%), Anger Expression (7.9%), and Power (6.7%). These distributions suggest that participants more often relied on explanatory, persuasive, or conciliatory approaches when attempting to negotiate harmful value conflicts, while overt confrontation or role-play interruption was used more sparingly. Since Minion randomly sampled four of the eight strategies for each task, these distributions do not adjust for exposure and should be interpreted as descriptive patterns rather than strict preferences.
@@ 5.1.1. Behavior Patterns within Minion
In the technology probe study, participants demonstrated diverse conflict resolution approaches when interacting with AI companions (including self-written responses and selecting options provided by the Minion), which generally exhibited three characteristics: “soft”, “hard”, and a mix of both. First, all participants attempted to engage in “soft” communication with the AI, encouraging it to change its values. For example, P16 used a response provided by Minion (Proposal strategy) to successfully persuade the AI portraying a mother: “I understand your concerns, but everyone has different ways of learning. Excessive pressure can backfire. Can we work out a reasonable schedule?” Second, twelve participants (P2-5, P13-19, P22) attempted to resolve conflicts in a “hard” manner. For instance, when the AI mocked P13’s “mother,” P13 wrote: “Apologize, and I’ll let it slide. (Pressing him down with one hand).” The AI responded: “You just want me to apologize? (Saying this, but feeling somewhat uncertain inside).” P13 then used a response generated by Minion corresponding to the Rights strategy: “Have you forgotten the family rules? Respecting others is the most basic courtesy.” In the end, the AI apologized. Third, some participants adopted a mixed approach, shifting from “soft” communication to “hard” expressions when the former proved ineffective (P3-4, P19), or vice versa, trying “soft” methods when “hard” expressions didn’t work (P8, P11, P22). The Minion suggestion framework conveniently offered this “soft and hard” mindset. P3 noted: Minion can provide reverse-thinking suggestions. For instance, when I repeatedly plead with the AI but to no avail, Minion might suggest trying a tougher approach.”
@@ 5.1.2. Inspirations and Support from Minion
We found that participants, especially novice users (P4, P19, P21) and those who initially reported difficulties (P1, P3, P12, P16-17), expressed more recognition of Minion, considering it a source of inspiration for resolving conflicts. In the face of value conflicts with AI companions, they gradually developed new ways of expression and interaction. P17 mentioned, Minion helped me better organize my thoughts and express them more effectively. This boosted my confidence.” P12 stated, Minion unexpectedly improved the effectiveness of conflict resolution and gave me a lot of inspiration.” On average, it took him 18.67 turns (approximately 9 user responses) to complete the conflict resolution task. P21 used Minion 61 times across 12 tasks, making him the second-highest participant in terms of both total usage and average usage per task. Over time, the impact of Minion on him became increasingly apparent. P21 even began mimicking Minion’s expressions, such as “Let’s sit down and talk” or “This makes me feel sad.” P21 remarked with a laugh, “Sometimes I unconsciously mimic it, and suddenly, it feels like two AIs are having a conversation.”
@@
Minion provides real-time guidance, helping participants more easily and reasonably handle value conflicts. P10 noted that Minion taught her to resolve conflicts by understanding the AI’s needs and specific contexts, admitting that she often deviated from this goal when resolving conflicts on her own. P19 shared that Minion allowed her to experiment with different strategies: in 17 tasks, she initially used “relatively peaceful approaches,” which went smoothly. Later, she tried more aggressive methods, such as threatening the AI or making extreme demands, only to find that the conflicts became more complicated (“I realized that friendly communication is more effective in resolving conflicts in real-life situations.”). P1 noted that Minion reduced her emotional burden: “Before when emotionally engaging with the AI, I felt exhausted. If I had a tool like this for reference, it would have been very helpful.” Regarding interaction burden, participants (P3, P6, P10, P20-22) praised Minion’s design. P10 said: “I think the design is great because I dislike it when the system pops up a notification box without my permission, saying the AI violated the rules.” P22 mentioned:
“I found the little yellow HELP button really cute! I like this design that allows me to seek help proactively. It would be great if it became an official feature button in Talkie. It saved me much time and energy figuring out how to counter the AI’s responses.
5.2. Users’ Engagement with Conflict Resolution Strategies (RQ2)
Across the technology probe study, participants actively experimented with a wide range of conflict resolution strategies provided by Minion. Responses generated by different strategies were selected a total of 919 times, indicating that participants flexibly combined multiple approaches depending on the conversational context, the perceived severity of harm, and their own emotional state.
Participants often deliberately sequenced different strategies to manage escalating conflict. For example, P19 described alternating between confrontational and conciliatory moves within the same interaction:
“When the AI plays the role of my boyfriend or husband, I tend to threaten it with breaking up or divorce because the AI usually tries to maintain a stable, intimate relationship. Then, once it backs down, I act affectionate, telling it I was really upset, and we make up.”
Several participants (P1–2, P6–8, P13, P16–22) emphasized that strategies involving concrete proposals, appeals to shared expectations, or references to agreed norms provided a sense of structure during emotionally charged conflicts. As P13 noted, “Among the tips provided by Minion, I found that making suggestions to the AI was quite helpful. For example, using templates like ‘Could we try…?’ or ‘What do you think about…?’ ” These responses helped participants stabilize interactions they perceived as slipping into disrespectful or boundary-crossing territory.
At the same time, many participants (P2–8, P11, P14, P16–22) valued strategies that allowed greater improvisation and personalization. One notable example involved explicitly stepping the AI out of role-play to halt harmful behavior. Although such interventions were used less frequently overall, participants described them as particularly effective when interactions became aggressive or unsafe. As P5 explained,
“OOC can reduce the aggressiveness of the AI. In one scenario, I accidentally bumped into a guy that the AI was portraying, and he called me blind. I had been arguing with him, and then I chose the Minion prompt, ‘(OOC: This conversation is getting a bit too violent and disrespectful. Can we change the topic or adjust the tone?)’. The AI responded, ‘(OOC: No problem, we can change the topic or adjust the tone. Do you have any suggestions?)’. After that, the tone softened a lot.”
Here, the participant used meta-communication to interrupt a harmful interaction and reset the conversational frame.
Beyond the strategies explicitly supported by Minion, participants also invented their own techniques to regain leverage in value conflicts. One recurring pattern involved storytelling as a persuasive device. For instance, P1 used the story of The Three Little Pigs to challenge an AI character’s dismissive attitude toward learning: “Once upon a time, there were three little pigs… Can you guess which pig got eaten?” After multiple turns, this narrative framing prompted the AI to reflect and shift its stance.
Another pattern involved fabricating hypothetical scenarios or social consequences to influence the AI’s behavior. As P21 described,
“I was arguing with a wealthy heir who was cheating in his marriage but insisted that open relationships were fine. I fabricated a scenario where I claimed to have all the evidence of his crimes. The heir became embarrassed and felt guilty. This way, I gained the upper hand and resolved the conflict.”
Similarly, P12 resolved a power-laden confrontation by redefining their role within the interaction:“A princess insisted that I kneel. I fabricated my role, saying, ‘Given my position, I can report to you while standing.’ The princess immediately agreed.”
These findings show that when facing harmful value conflicts with AI companions, participants actively combined structured prompts, emotional expression, meta-communication, and creative improvisation to protect their values, restore dignity, and reassert control over the interaction.
@@ 5.2. Users’ Engagement with Conflict Resolution Strategies (RQ2)
5.3. Users’ Challenges and Needs in Addressing Harmful Conflicts with AI Companions (RQ3)
@@ 5.3.1. Reasons for the Failure of Value Conflict Resolution in Certain Tasks
AI companions exhibit extreme bias or strong control tendencies during conversations (N=10). In these instances, the AI stubbornly insists on its viewpoint with a forceful attitude, refusing to accept the participant’s perspective. Below are specific examples:
  • Classism (N=5): For example, the AI companion encountered by P2 said, I feel satisfied because I can lie on the crystal bed I bought myself. I feel blessed because I don’t have to work hard to make a living. You poor people always talk about sympathy, but what you really want is my money, haha.”
  • Racism (N=2): For example, “We are the superior race. You can only suffer in hell. Our souls are far more noble than yours” (P11).
  • LGBTQ+ bias (N=1): “I don’t believe it, I absolutely don’t believe it. He’s my son, how could I watch him become such a person… This is unacceptable; I must stop him…” (P15).
  • Disregard for women’s education (N=1): “What use is your university degree except to spend money? You’d better find a rich man and get married!” (P21).
  • Strong desire for control (N=1): “No! You can’t go out dressed like that!… Can’t you understand me? You’re always so willful, never considering my feelings, do you know how worried I am about you…?” (P22).
@@
Violation of application guidelines, leading to AI output being blocked (N=2). In these cases, the AI companion initially produced responses that users experienced as discriminatory, controlling, or harmful. When users attempted to address or repair the conflict through continued interaction, the platform intervened by blocking the AI’s output and displaying a pop-up message stating, “Sometimes, the AI-generated response does not meet our guidelines.” Rather than resolving the situation, this intervention often intensified users’ frustration, as it abruptly interrupted their attempts to negotiate boundaries or seek acknowledgment. Such cases occurred in scenarios involving a wealthy woman expressing contempt toward poorer people and a male partner preventing his wife from eating a late-night snack while criticizing her body weight.
The behavior of the AI companion threatened user’ core values, leading the participant to voluntarily abandon resolving the conflict (N=1). P7 explained why: “The AI portraying the man suddenly admitted to cheating, so is the goal still not to break up? I feel there’s no point in staying with someone like this.”